Find the z – transform of z(cos tita) and z(sin tita)


1
Expert's answer
2020-09-16T17:46:19-0400

since

"cos\theta=real \, part\, of \,e^{i\theta}"

sin"\theta" =Imaginary part of "e^{i\theta}"


let calculate z-transform of "e^{i\theta}"

Z{"e^{i\theta}" }= "\frac{z}{z-e^{i\theta}}"


="\frac{z}{z-cos\theta-isin\theta}"

="\frac{z}{z-cos\theta-isin\theta}\times \frac{z-cos\theta+isin\theta}{z-cos\theta+isin\theta}"


="\frac{z(z-cos\theta)+izsin\theta}{(z-cos\theta)^2-i^2sin^2\theta}"


="\frac{z(z-cos\theta)+izsin\theta}{(z^2+cos^2\theta-2zcos\theta+1sin^2\theta}"


="\frac{z(z-cos\theta)+izsin\theta}{(z^2+1-2zcos\theta)}"


on separating real and imaginary parts we get,


we get

"Z( {cos\theta}) =\frac{z(z-cos\theta)}{(z^2+1-2zcos\theta)}"


"Z(sin\theta)=\frac{zsin\theta}{(z^2+1-2zcos\theta)}"